Can someone please help me with this problem. Thank you
f(x)=x²+7 and g(x)=√x-2 and find (g0f)(x)
Thanks again

Answer by fractalier(6550)

(g o f)(x) means g(f(x))...
Here that means
g(f(x)) = g(x^2 + 7) = sqrt(x^2 + 7 - 2) =
sqrt(x^2 + 5)